Hello. Please solve the following below. Show solution.


1.      How many significant digits in 0.01850 – 103?

2.      How many significant digits in 219900?

3.Calculate 66.666*33.333 and give the 4 significant digit answer.


1
Expert's answer
2021-10-21T14:33:12-0400

1.

"0.01850 \u2013 10^3=-999.9815"

The Significant Figures are "\begin{matrix}
 \ 9 & 9 & 9 & 9 & 8 & 1 & 5 \\
\end{matrix}"

There are seven Significant Figures.


2. "219900"

The Significant Figures are "\begin{matrix}
 \ 2 & 1 & 9 & 9 \\
\end{matrix}"

There are four Significant Figures.


3.


"66.666\cdot33.333=2222.177778"

Give the 4 significant digit answer


"2222"